### v1

@@ 4 Data and Descriptive Statistics
Starting with the Cohesion Fund data, we note that they are generally available for the 1993-2022 period. [^7] Although the Cohesion Fund was officially launched in 1994, there were some payments in advance towards the eligible countries in 1993. However, the vast majority of regions became eligible for the CF payments in 2000 (as already mentioned, prior to 2000, only four countries received Cohesion Fund payments). Given that the data are provided in nominal terms, we convert them to 2015 prices using the regional GVA deflator. We note here that prior to using the data for the empirical analysis, we had to account for two issues: first, the fact that the data are not reported under a single NUTS2 classification; and, second, the issue of the time of the recording of Cohesion Fund expenditures. Further details regarding these issues, along with our approach, are available in Appendix B.
@@
Panel (a) of Figure 1 depicts the evolution of the CF payments over time for all recipient countries, while panel (b) of the Figure depicts the share of disbursements received by the eligible countries. As is evident from the Figure, during the 2007-2013 period there was a significant increase in the CF payments as a means of accommodating the accession of the newest Member States. These increased payments continued up to 2015, due to the well-known delays in the disbursement of EU funds (for more details on this, see Dicharry (2023)). In total, for the period covered by our sample, €165.15 billion were disbursed to the recipient countries, with Poland and Spain receiving around 27% and 17% of total payments, respectively.
Figure 1: Cohesion Fund payments
(a) Yearly Cohesion Fund payments
(b) Share of Cohesion Fund payments
@@
Notes: Panel (a) depicts the total expenditures for the Cohesion Fund by year, while panel (b) presents the average shares of expenditures received by the 17 recipient countries. Note that not all countries received Cohesion Fund payments throughout the sample, e.g., Spain was ineligible for the 2014-2020 period.
@@
Table 3 provides a disaggregation of the CF disbursements by beneficiary, along with some additional data on the relative importance of these payments in terms of GVA and public investment. The Table covers the 2000-2022 period, which includes the last three programming periods, in order to ensure maximum coverage in terms of recipient countries. We note here that Ireland is excluded from the analysis for two reasons: firstly, Ireland was eligible for the Cohesion Fund for only one programming period (2000-2006). Secondly, due to the well-documented issues with the Irish National Accounts stemming from the distortionary impact of large multinational enterprises operating within the country it would be increasingly difficult to disentangle the effects of the Cohesion Fund on economic performance. Overall, it is evident from the Table that Poland has received the bulk of the payments over the last 23 years, with a total amount that is more than twice in magnitude compared to the one received by Spain, which ranks second. It is worth highlighting here that Spain was not eligible for funding from the CF during the last programming period (2014-2020), implying that the country benefited significantly from these disbursements within a narrower time frame.
@@
Turning to the relative size of the CF payments in terms of GVA, we observe that across beneficiaries the CF ranges from 0.08 to 0.68%, with a median value of 0.39%. As such, the contribution of the EU’s CF transfers constitute a significant boost of liquidity for the implementation of investment projects. This is further exemplified in the last column of the Table, which depicts the CF transfers as a share of public investment. We observe that, with a median value of 11.7%, the CF transfers emerge as a significant source of additional funds for the financing of investment, in a complementary fashion to public investment. In particular, in Eastern countries like Lithuania, Latvia, Hungary, Slovakia and Poland, the CF disbursements amounted to at least 15% of public investment.
@@
Country Regions CF amount CF(GVA) CF (Gov. Investment)
@@
Our source for macro-related data is ARDECO. We obtain data for real Gross Value Added (GVA), at constant 2015 prices (series code: SOVGE) and for the average annual population at the regional level (series code: SNPTD) in order to construct the corresponding per capita measure.
Our dataset originally included data from 27 European countries and $N=242$ regions in the 2021 NUTS2 classification for a maximum of $T=43$. The treated sample covers 96 EU NUTS2 regions [^8] We exclude from the sample of treated regions the so-called outermost regions of Portugal, namely Azores and Madeira, and Spain, namely the Canary Islands, along with the autonomous cities of Ceuta and Melilla. over the 1993-2022 period for a total of 2297 observations and an average of 23.9 years under treatment. These regions belong to 16 countries, the majority of which are the newest Member States of the EU, i.e. the Central and Eastern European countries that entered the Union starting with the 2004 enlargement. Only 18 regions (located in Greece and Portugal) are under treatment for the whole sample period, with 78 regions being treated for up to 23 years.
@@ 5.1 Notation and preliminaries
We work under the framework of potential outcomes (see Holland (1986)) to assess the impact of the policy intervention implied by the CF program by relying on observational time-series. For each region and at each time period two potential outcomes are considered; the outcome that region $i$ would have if it was not treated at time $t$, denoted by $y_{it}(0)$ and typically referred as the potential untreated outcome, and the outcome that the $i$-th region would have at time $t$ if was treated, namely the potential treated outcome denoted by $y_{it}(1)$. Notice that only one of these outcomes is observed, while the other is known as the counterfactual outcome. Specifically, for a treated region $i\in\mathcal{N}$ at time $t>T_{i}$ we observe $y_{it}=y_{it}(1)$ whereas for $t\leq T_{i}$ as well as for $i\notin\mathcal{N}$ and for each $t=1,\ldots,T$ we observe $y_{it}=y_{it}(0)$. The causal effect $\tau_{it}$ of the policy intervention is defined in terms of the difference between the observed outcome and the one that region $i$ would have if it was not treated at time $t$, i.e. the counterfactual, so that
@@
Finally, let $D_{it}$ denote the treatment indicator, $D_{it}=1$ if region $i$ received the intervention at time $t$ and $D_{it}=0$ otherwise.
@@ 5.2 Causal inference by using matrix completion
To assess the effect of the policy intervention implied by the CF program we follow recent directions in the literature and particularly the factor augmented approach (Xu, 2023) where unobserved confounders are modeled through interactive fixed effects (Bai, 2009b). Specifically, we employ the matrix completion method (Athey et al., 2021, Borusyak et al., 2024), considering the unobserved potential untreated outcomes $y_{it}(0)$, $i\in\mathcal{N}$ and $t>T_{i}$, as missing values to be estimated as $\widehat{y}_{it}(0)$. This allows us to compute the causal effect $\tau_{it}$ as
@@
In particular, we assume for each $i=1,\ldots,N$ and $t=1,\ldots,T$ the following factor model for both observed and potential untreated outcomes
$y_{it}$ $=\tau_{it}D_{it}+\lambda_{i}^{\top}f_{t}+\epsilon_{it}$ (3)
$=\lambda_{i}^{\top}f_{t}+\epsilon_{it}$
where $f_{t}$ is a $K\times 1$ vector with unobserved factors, $\lambda_{i}$ is $K\times 1$ vector with factor loadings and $\epsilon_{it}$ is a random variable that accounts for measurement error and is assumed to follow a sub-Gaussian distribution; the second line in (3) emphasizes that for potential untreated outcomes we have $D_{it}=0$ for all $i=1,\ldots,N$ and $t=1,\ldots,T$, i.e,. (3) can also be interpreted as our assumption for the data generation mechanism of untreated outcomes. Moreover, the model in (3) is a generalization of the Two Way Fixed Effects (henceforth, TWFE) model which is the standard tool for estimation of causal effects under the DiD approach; see Appendix A for more details. Crucially, causal estimation using the TWFE model aims to estimate the parameter $\tau=\mathbb{E}[\tau_{it}|D_{it}=1]$ which is typically known as the average treatment effect on the treated (ATT) while disregarding the estimation of the unit- and time- specific effects $\tau_{it}$. Additionally, the estimation of $\tau$ by using the TWFE modelling approach relies on the following set of strong assumptions: (i) additive separability which states that any unobserved heterogeneity can be decomposed into additive unit- and time- specific fixed effects, (ii) the parallel trends assumption, which states that the treated and control outcomes would follow the same paths in the absence of the treatment, (iii) strict exogeneity of treatment, which rules out the possibility of the treatment being assigned based on future outcomes, (iv) no dynamic treatment effects, i.e., treatment effects are assumed to be homogeneous and constant over time and (v) the stable unit treatment value assumption (SUTVA) which states that the potential outcomes of any given unit are not affected by the treatment status of other units.
Importantly, the model specification described by (3) mitigates or rules out most of the aforementioned assumptions. First, the parallel trends assumption is relaxed since unobserved time-varying confounders that cause its failure are modelled by the interactive fixed effects term $\lambda_{i}^{\top}f_{t}$. In fact, the latent factors $f_{t}$ can be understood as time-varying parameters that model unobserved confounders which are common across the EU regions but exert a heterogeneous impact on each region; this region-specific effect of the unobserved confounders is the parameter $\lambda_{i}$, also known as the factor-loading parameter. Furthermore, the interactive fixed effects term rules out the additive separability accounting for richer cross-sectional dependence. Notably, this is particularly useful in regional policy evaluation, where units considered to be in close proximity (either geographical or based on economic considerations, e.g. due to trade) may respond differently to unobserved shocks. Furthermore, as already noted, our considered approach results in the estimation of heterogeneous causal effects $\tau_{it}$ across time and regions, with the ATT for each time $t$ being estimated as
$\widehat{ATT}_{t}=\frac{1}{N_{tr}}\sum_{i\in\mathcal{N}}\widehat{\tau}_{it}.$ (4)
Overall, the identification of causal effects provided by the factor modelling approach in (3) relies on the following two assumptions: (i) strict exogeneity of treatment, i.e., treatment assignment is assumed to be uncorrelated with the idiosyncratic error after controlling for the latent factors and thus rules out endogenous treatments (ii) the potential outcomes of one unit should not be affected by the treatment status of another unit. Notice, that (ii) is a relaxed form of SUTVA since the interactive effects in (3) can capture unobserved common shocks, such as sectoral linkages, trade flows, or institutional similarities that may lead to correlated outcomes.
@@ 5.2.1 Estimation of causal effects
To estimate $y_{it}(0)$, for $i\in\mathcal{N}$ and $t>T_{i}$, we fit the model in (3) by discarding all the post-intervention data. In particular, following Athey et al. (2021) we consider the partially observed outcome matrix $Y\in\mathbb{R}^{N\times T}$ in which
$y_{it}=\begin{cases}y_{it}(0),&\text{if $i\notin\mathcal{N}$ and $t=1,\ldots,T$},\\ y_{it}(0),&\text{if $i\in\mathcal{N}$ and $t\leq T_{i}$},\\ \text{missing},&\text{$i\in\mathcal{N}$ and $t>T_{i}$}.\end{cases}$
Then, the model in (3) implies that
$Y=L+E,$ (5)
where $L$ is a $N\times T$ low-rank matrix representing the underlying structure of the data with elements $L_{it}=\lambda_{i}^{\top}f_{t}$ and $E$ is also a $N\times T$ matrix with elements $\epsilon_{it}$. Clearly, from (5) we have both for the observed and the missing entries in $Y$ that
$y_{it}=L_{it}+\epsilon_{it},$
The goal is to estimate the entries $L_{it}$ that correspond to the missing entries in $Y$. To achieve this we employ the matrix completion with nuclear-norm regularization method developed recently by Athey et al. (2021). In particular, given the observed entries in $Y$ indexed by $\mathcal{O}\subset\{1,\dots,N\}\times\{1,\dots,T\}$, the estimation of $L$ is formulated as a convex optimization problem of the form
$\hat{L}=\arg\min_{L}\|P_{\mathcal{O}}(Y-L)\|_{F}^{2}+\lambda\|L\|_{*},$ (6)
where $P_{\mathcal{O}}((Y-L))$ denotes the projection operator that retains entries in $\mathcal{O}$ and sets others to zero, $\|P_{\mathcal{O}}(Y-L)\|_{F}$ is the Frobenius norm of the matrix $P_{\mathcal{O}}(Y-L)$, $\|L\|_{*}$ is the nuclear norm (sum of singular values), which promotes a low-rank solution and $\lambda$ is a tuning parameter controlling the trade-off between fit and complexity; see Appendix A for the definitions of the Frobenius and nuclear norms as well as for more details on solving the minimization problem in (6). Finally, having at hand the estimators $\hat{L}_{it}$ we estimate the causal effects for the treated units, i.e., for $i\in\mathcal{N}$ and $t\geq T_{i}$, as
$\hat{\tau}_{it}=y_{it}-\hat{y}_{it}(0)=y_{it}-\hat{L}_{it}.$

### v2

@@ 4 Data and Descriptive Statistics
Starting with the Cohesion Fund data, we note that they are generally available for the 1993-2022 period. [^11] Although the Cohesion Fund was officially launched in 1994, there were some payments in advance towards the eligible countries in 1993. However, the vast majority of regions became eligible for the CF payments in 2000 (as already mentioned, prior to 2000, only four countries received Cohesion Fund payments). Given that the data are provided in nominal terms, we convert them to 2015 prices using the regional GVA deflator. We note here that prior to using the data for the empirical analysis, we had to account for two issues: first, the fact that the data are not reported under a single NUTS2 classification; and, second, the issue of the time of the recording of Cohesion Fund expenditures. Further details on how we deal with these issues are available in Appendix C.
@@
Figure 2 depicts the evolution of the CF payments over time for all recipient countries, while Figure 2 depicts the share of disbursements received by the eligible countries. As is evident from the Figure, during the 2007-2013 period there was a significant increase in the CF payments as a means of accommodating the accession of the newest Member States. These increased payments continued up to 2015, due to the well-known delays in the disbursement of EU funds (for more details on this, see Dicharry (2023)). In total, for the period covered by our sample, €165.15 billion were disbursed to the recipient countries, with Poland and Spain receiving around 27% and 17% of total payments, respectively.
Figure 1: Yearly Cohesion Fund payments
Figure 2: Share of Cohesion Fund payments
@@
Note: Figure 2 depicts the total expenditures for the Cohesion Fund by year, while Figure 2) presents the average shares of expenditures received by the 17 recipient countries. Note that not all countries received Cohesion Fund payments throughout the sample, e.g., Spain was ineligible for the 2014-2020 period.
@@
Table 2 provides a disaggregation of the CF disbursements by beneficiary, along with some additional data on the relative importance of these payments in terms of GVA and public investment. The Table covers the 2000-2022 period, which includes the last three programming periods, in order to ensure maximum coverage in terms of recipient countries. We note here that Ireland is excluded from the analysis for two reasons: firstly, Ireland was eligible for the Cohesion Fund for the first two programming periods (1994-1999 and 2000-2006). Secondly, due to the well-documented issues with the Irish National Accounts stemming from the distortionary impact of large multinational enterprises operating within the country it would be increasingly difficult to disentangle the effects of the Cohesion Fund on economic performance. [^12] For the Ireland’s distorted national accounts see Kostarakos et al. (2024) Overall, it is evident from the Table that Poland has received the bulk of the payments over the last 23 years, with a total amount that is more than twice in magnitude compared to the one received by Spain, which ranks second. It is worth highlighting here that Spain was not eligible for funding from the CF during the last programming period (2014-2020).
@@
Turning to the relative size of the CF payments in terms of GVA, we observe that across beneficiaries the CF ranges from 0.12 to 0.68%, with a median value of 0.39%. As such, the contribution of the EU’s CF transfers constitute a significant boost of liquidity for the implementation of investment projects. This is further exemplified in the last column of the Table, which depicts the CF transfers as a share of public investment. We observe that, with a median value of 11.7%, the CF transfers emerge as a significant source of additional funds for the financing of investment, in a complementary fashion to public investment. In particular, in Eastern countries like Lithuania, Latvia, Hungary, Slovakia and Poland, the CF disbursements amounted to at least 15% of public investment.
@@
Country Regions CF amount CF(% GVA) CF (% Gov. Investment)
@@
Our source for macro-related data is ARDECO. We obtain data for real Gross Value Added –GVA– (series code: SOVGE) and Gross Fixed Capital Formation –GFCF– (series code: ) at constant 2015 prices. Moreover, we use data for the average annual population at the regional level (series code: SNPTD) and total employment (series code: SNETD). We log-transform the outcome variables, that is GVA per capita and the rate of investment. As such, the empirical results will be interpreted as the percentage change in each variable following the disbursement of of the Cohesion Fund transfers (i.e. after the treatment).
Our initial dataset consists of an unbalanced panel of 242 EU NUTS2 regions belonging to 27 EU Member States — based on the 2021 classification — and spans the 1985–2022 period. We exclude from the sample the so-called outermost regions of France, Portugal, and Spain, [^13] Specifically, we drop from the sample five French outermost regions, the Azores and Madeira from Portugal, and from Spain the Canary Islands along with the autonomous cities of Ceuta and Melilla. while we also exclude Ireland for the reasons mentioned above. As such, we end up with a sample of 228 EU NUTS2 regions and 8,249 observations. The sample of treated regions consists of 96 EU NUTS2 regions over the 1993–2022 period, amounting to a total of 2,297 observations and an average of 23.9 years under treatment. These regions belong to 16 countries, the majority of which are the newest Member States of the EU, i.e. the Central and Eastern European countries that entered the Union starting with the 2004 enlargement. Importantly, no region is treated before 1993 — the first year in which Cohesion Fund support becomes available. From that point onward, only 18 regions (located in Greece and Portugal) receive funding continuously for the entire treatment period (1993–2022), while the remaining 78 treated regions receive support for up to 23 years.
@@ 5.1 Notation and preliminaries
We work under the framework of potential outcomes (see Holland (1986)) to assess the impact of the policy intervention implied by the CF program by relying on observational time-series. For each region and at each time period two potential outcomes are considered; the outcome that region $i$ would have if it was not treated at time $t$, denoted by $y_{it}(0)$ and typically referred as the potential untreated outcome, and the outcome that the $i$-th region would have at time $t$ if it was treated, namely the potential treated outcome denoted by $y_{it}(1)$. Notice that only one of these outcomes is observed, while the other is known as the counterfactual outcome. Specifically, for a treated region $i\in\mathcal{N}$ at time $t>T_{i}$ we observe $y_{it}=y_{it}(1)$ whereas for $t\leq T_{i}$ as well as for $i\notin\mathcal{N}$ and for each $t=1,\ldots,T$ we observe $y_{it}=y_{it}(0)$. The causal effect $\tau_{it}$ of the policy intervention is defined in terms of the difference between the observed outcome and the one that region $i$ would have if it was not treated at time $t$, i.e., the counterfactual, so that for treated regions and $t>T_{i}$:
@@
Finally, let $D_{it}$ denote the treatment indicator, $D_{it}=1$ if region $i$ has received the intervention at time $t$ and $D_{it}=0$ otherwise.
@@ 5.2 Causal inference by using matrix completion
To assess the effect of the policy intervention implied by the CF program, we follow recent directions in the literature and particularly the factor-augmented approach (Xu, 2023), where unobserved confounders are modeled through interactive fixed effects (Bai, 2009b). Specifically, we employ the matrix completion method (Athey et al., 2021, Borusyak et al., 2024), considering the unobserved potential untreated outcomes $y_{it}(0)$, $i\in\mathcal{N}$ and $t>T_{i}$, as missing values to be estimated as $\widehat{y}_{it}(0)$. This allows us to estimate the causal effect $\tau_{it}$ as
@@
In particular, we assume for each $i=1,\ldots,N$ and $t=1,\ldots,T$ the following factor model for both observed and potential untreated outcomes becomes
$y_{it}$ $=\tau_{it}D_{it}+X_{it}^{\top}\beta+\lambda_{i}^{\top}f_{t}+\epsilon_{it}$ (3)
$=X_{it}^{\top}\beta+\lambda_{i}^{\top}f_{t}+\epsilon_{it}$
where $X_{it}$ is a $p\times 1$ vector of observed covariates and $\beta$ the corresponding coefficient vector. [^15] Following Gobillon and Magnac (2016) and Xu (2017), among others, we assume that $\beta$ is constant across regions and time. This is justified by the use of interactive fixed effects, which flexibly capture complex region-time heterogeneity. Allowing $\beta$ to vary across $i$ and $t$ introduces estimation complexity and risks confounding observed and unobserved components; see also Athey et al. (2021). The covariates $X_{it}$ account for observed sources of heterogeneity and improve identification under the standard conditional independence assumption
$\{Y_{it}(1),Y_{it}(0)\}\perp D_{it}\mid X_{it},\lambda_{i},f_{t}.$
In practice, this condition is typically satisfied when $X_{it}$ captures pre-treatment or predetermined characteristics. Conditioning on post-treatment variables is not required for identification and can change the estimand by blocking causal channels. The term $f_{t}$ is a $K\times 1$ vector of unobserved common factors, and $\lambda_{i}$ is a $K\times 1$ vector of region-specific loadings; together, they flexibly capture unobserved time-varying heterogeneity and cross-sectional dependence. The second line of (3) emphasizes that for potential untreated outcomes we have $D_{it}=0$ for all $i$ and $t$, i.e., (3) also describes the data-generating process of untreated outcomes.
Moreover, the model in (3) generalizes the standard Two-Way Fixed Effects (TWFE) specification, which is widely used for causal estimation under the Difference-in-Differences (DiD) framework; see Appendix B for more details. Under its identifying assumptions, the TWFE estimator recovers the parameter $\tau=\mathbb{E}[\tau_{it}\mid D_{it}=1,X_{it}]$, the average treatment effect on the treated (ATT), but it cannot capture heterogeneous, unit- and time-specific effects $\tau_{it}$. Estimation based on the TWFE model relies on several strong assumptions: (i) additive separability, requiring unobserved heterogeneity to be captured by additive unit and time fixed effects; (ii) parallel trends, which assumes treated and control units would evolve similarly absent treatment; (iii) no anticipation, excluding selection on future outcomes; and (iv) homogeneous treatment effects, implying constant effects across units and time.
The specification in (3) mitigates reliance on most of these assumptions. First, the parallel trends condition is relaxed because unobserved time-varying confounders that could violate it are captured by the interactive fixed effects term $\lambda_{i}^{\top}f_{t}$. The inclusion of observed covariates $X_{it}$ further helps account for observable differences between treated and control units, improving balance and reducing residual variance, provided these covariates are predetermined and not themselves affected by treatment. The latent factors $f_{t}$ represent unobserved time-varying shocks common to all regions, while $\lambda_{i}$ allows these shocks to affect regions heterogeneously. This richer structure captures complex forms of cross-sectional dependence that cannot be addressed by additive fixed effects alone, thereby mitigating the additive separability restriction. Finally, allowing treatment effects to vary across both units and time directly relaxes the homogeneous treatment effects assumption — a particularly important feature in the context of regional policy evaluation, where differences in economic structure, institutional capacity, or spatial linkages are expected to generate heterogeneous responses to policy interventions.
Importantly, our approach enables the estimation of heterogeneous causal effects $\tau_{it}$ across units and time. We summarize these effects via a time-specific $ATT_{t}$ (commonly called an event-study estimate), as
$\widehat{ATT}_{t}=\frac{1}{|N_{tr,t}|}\sum_{i\in\mathcal{N}}\widehat{\tau}_{it},$ (4)
where $N_{tr,t}=\{i:T_{i}\leq t\}$ is the set of units treated by time $t$. Overall, the identification of causal effects provided by the factor model in (3) relies on two main assumptions: (i) strict exogeneity of treatment, i.e., treatment is uncorrelated with the idiosyncratic error term after controlling for both the latent factors and the observed covariates $X_{it}$, thus ruling out endogenous selection; and (ii) no interference between units, in the sense that the potential outcomes of a given unit should not depend directly on the treatment assignment of other units. This second condition is partly relaxed in practice by the presence of interactive fixed effects, which can accommodate correlated unobserved shocks such as trade linkages or coordinated policy responses.  [^16] While interactive fixed effects can accommodate latent cross-sectional dependence, they do not fully address violations of the no-interference assumption. Explicit modeling of spillovers in panel causal inference remains challenging and requires additional structure or detailed network information often unavailable in regional policy settings. We therefore follow the state-of-the-art approach for imputing counterfactual outcomes by combining interactive fixed effects with matrix completion, building on the foundational work of Gobillon and Magnac (2016) and recent advances in low-rank causal imputation (e.g., Athey et al., 2021).
@@ 5.2.1 Estimation of causal effects
To operationalize the model in (3), we cast the estimation of counterfactual untreated outcomes $y_{it}(0)$ as a matrix completion problem. As shown in Athey et al. (2021), the interactive fixed effects structure can be equivalently expressed at the matrix level by decomposing the full outcome matrix $Y$ into a low-rank matrix capturing latent effects, a covariate component, and noise. Specifically, the estimation model becomes:
$Y=L+X\beta+E,$ (5)
which corresponds to a matrix reformulation of (3), where the latent factor term $\lambda_{i}^{\top}f_{t}$ across $i$ and $t$ is compactly represented by the low-rank matrix $L$; $L_{it}=\lambda_{i}^{\top}f_{t}$. The matrix $X$ stacks the covariates $X_{it}$ appropriately across units and time. Thus, (5) can be viewed as the global form of the unit-time model in (3), and provides a natural starting point for matrix completion techniques to estimate the missing entries corresponding to $y_{it}(0)$.
In particular, we observe a subset of entries in $Y$, indexed by $\mathcal{O}\subset\{1,\ldots,N\}\times\{1,\ldots,T\}$, and estimate $(L,\beta)$ by solving the following regularized convex optimization problem
$(\hat{L},\hat{\beta})=\arg\min_{L,\beta}\left\{\|P_{\mathcal{O}}(Y-L-X\beta)\|_{F}^{2}+\lambda\|L\|_{*}\right\},$ (6)
where $P_{\mathcal{O}}(\cdot)$ is the projection operator that retains entries in $\mathcal{O}$ and zeros out the rest $\|\cdot\|_{F}$ denotes the Frobenius norm, $\|\cdot\|_{*}$ denotes the nuclear norm (sum of singular values), $\lambda>0$ is a tuning parameter that controls the rank of $\hat{L}$.
The optimization problem in (6) is solved via an iterative procedure, alternating between (i) estimating $\hat{\beta}$ using least squares on the residuals YLY-L over observed entries and (ii) estimating $\hat{L}$ using a matrix completion algorithm (e.g., soft-impute) on the residuals $Y-X\hat{\beta}$. This procedure continues until convergence ; see in the Appendix XX for more details. Once $\hat{L}$ and $\hat{\beta}$ are obtained, the counterfactual untreated outcomes are imputed as
$\hat{y}_{it}(0)=\hat{L}_{it}+X_{it}^{\top}\hat{\beta},\quad\text{for }i\in\mathcal{N},\,t>T_{i}.$
Hence, the estimated treatment effects for treated units in post-treatment periods are
$\hat{\tau}_{it}=y_{it}-\hat{y}_{it}(0)=y_{it}-\hat{L}_{it}-X_{it}^{\top}\hat{\beta}$.
This approach separates latent structure from observed heterogeneity, improving both interpretability and predictive accuracy when covariates are informative.